OpenStudy (gudetamas):

which functions graph has a vertex at (3,5) and contains the point (5,13) y = 1/10(x+3)^2 - 5 y= 1/10(x-3)^2 - 5 y= 2(x-3)^2 + 5 y= 2(x+3)^2 + 5

OpenStudy (knawk):

C. y= 2(x-3)^2 + 5

OpenStudy (gudetamas):

@knwak could u show me the steps?

OpenStudy (gudetamas):

@Knawk

OpenStudy (tkhunny):

The all look like this: y = a(x-h)^2 + k You are given the vertex. That's (h,k) = (3,5) y = a(x-3)^2 + 5 All we need it 'a'. You are given a point. That't (x,y) = (5,13) 13 = a(5-3)^2 + 5 Solve for a. Show us your best algebra.

OpenStudy (gudetamas):

a would it be 2 ? @tkhunny

OpenStudy (tkhunny):

I don't know. You tell me - with confidence.